OpenStudy (anonymous):

What is the inverse of g of x equals the quantity x plus 8 divided by 3?

OpenStudy (anonymous):

does your equation look like \[g(x) = \frac{ X + 8 }{ 3 }\] ? to make it easier i replace the g(x) with y, then switch the x and y values then solve for y again so it would look like this: \[X = \frac{ y + 8 }{ 3 }\] then you should solve for y.

OpenStudy (anonymous):

yes it does look like the first one but I am having trouble solving for y

OpenStudy (anonymous):

So the first step, in my opinion would be to get the 3 to the other side. To get rid of the 3 you should multiply both sides by 3.

OpenStudy (anonymous):

so it would look like this, \[3x = y + 8\]

OpenStudy (anonymous):

should my final answer be x=3x-8?

OpenStudy (anonymous):

yeah, except it would be Y = 3x - 8 is the inverse of g(x)

OpenStudy (anonymous):

is this textbook work?

OpenStudy (anonymous):

no online school homework

OpenStudy (anonymous):

is there a way for you to check your answers?

OpenStudy (anonymous):

its multiple choice if that helps? but other than that no

OpenStudy (anonymous):

ah, well goodluck. Did that make sense to you?

OpenStudy (anonymous):

yes, thank you!

OpenStudy (anonymous):

No problem